OpenStudy (anonymous):

Find the values of each of the other trigonometric functions of u: (1.) tan u= -1 sec u= -2 (2.) csc u= -√ ̅ -2 and tan u <0

OpenStudy (anonymous):

Oh! My apologies. Those were just to differentiate two separate problems.

OpenStudy (mathmale):

Would you mind presenting these problems just one at a time? It takes more imagination on the part of the reader tha should be necessary to figure out what you are trying to say. I'll take a guess and ask whether you mean the following or not: If: tan u= -1, and sec u= -2, find the values of the other four trig problems. It's very important that you present your math problems in such a way that they are clear, not ambiguous. Notice how I have written tan u = -1 on one line, and sec u = -2 on the next line, to keep them separated. Have you begun thinking about how YOU would solve this problem? How are the secant and the tangent functions defined? What would your angle u look like, if tan u = -1 and sec u = -2?

OpenStudy (anonymous):

I'm sorry, this is my first time on the website. I'll make sure to present my questions in a more understandable manner in the future. I understand these types of problems but when it gets to questions with just one number (i.e. -1). For example, I know that if sin u= 5/13 that the opposite side would be 5 and the hypotenuse would be 13. From there I would be able to solve the rest, but with one number I become lost. If tan were -1 then the angle would either have to be in the first or fourth quadrant, but since it is negative, I know that it would be in the fourth.

OpenStudy (mathmale):

"If tan were -1 then the angle would either have to be in the first or fourth quadrant." How so? All of the trig functions are positive if the angle in question is in the 1st quadrant. Thus, if the tangent were -1 (which is actually opp/hyp = -1/1), then the angle u would not be in Q1 or Q3; it would have to be in either Q2 or Q4. If sec u = -2, think of this as sec u = -2/1. In other words, according to the definition of the secant function, sec u = hyp / adj = -2 / 1. This is legitimate; while the hyp is always positive, the adj. side can be either pos. or neg. So you have 2 requirements for the location of angle u: tan u is -1/1 and sec u is -2/1. In which quadrant will you find this angle, and why?

OpenStudy (anonymous):

Sorry, I got mixed up with angle rotations and finding angles of inverse trigonometric functions mathematically. Would it be in the second quadrant since secant is hypotenuse over adjacent and the hypotenuse would not be negative? It wouldn't be in the fourth quadrant because in the fourth quadrant, the y is negative and cosine, the opposite of secant, is x.